OpenStudy (anonymous):

HI, what is 0.08333333333 as a fraction? Could I have the regular answer and the simplified one? (if it can be simplified? ) Also the question is 11/16 times 4/3.

OpenStudy (anonymous):

its 4/33 not 4/3 sorry.

OpenStudy (akashdeepdeb):

Let 0.83bar = p 8.3bar = 10p 83.3bar = 100p 100p - 10p = 75 p = 75/90 = 25/30 = 5/6 = 0.83bar Understood? :)

OpenStudy (anonymous):

yes, thank you!

OpenStudy (akashdeepdeb):

\(\frac{11}{16} * \frac{4}{33}\) 1/4 * 1/3 = 1/12 Understood? :)

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

this is a really cool site...can you actually see my reply before I send it????

OpenStudy (akashdeepdeb):

Hahaha..No I cannot. And yes this is like an educational facebook! :D

OpenStudy (anonymous):

lol right? XD

OpenStudy (akashdeepdeb):

:)

OpenStudy (asnaseer):

@music_rocks_101 -- Welcome to OpenStudy The usual workflow on this site is: 1. You post a question. 2. One or more people come to your question to help you out. 3. You select the one that best answered your question by clicking on the blue Best Response button next to any one of their replies. 4. You close the question by clicking on the Close button then the whole process starts again. So you could start by clicking the Best Response against @AkashdeepDeb - I'm sure it will be appreciated :)

OpenStudy (anonymous):

thanks! @asnaseer

OpenStudy (asnaseer):

yw :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."